Omg I am so sorry but I am seriously so confused!
x+3y=6
how can I create a table of values for this?
y=6-x? I am honestly so confused
I got for x+y=6
y=6-x
And I managed to do the table and graphy and everything
I'm only 13!
Thank you so much

Answer by Alan3354(69443)

x+3y=6
how can I create a table of values for this?
-----------------
Solve for y
---
x+3y=6
3y = 6-x
y = 2 - x/3
--
Pick values for x and find y
x = 0, y = 2
x = 3, y = 1
x = 6, y = 0
etc